### v3

# Ultra-High-Energy Cosmic Rays from Neutrino-Emitting Tidal Disruption Events

Abstract
We revisit the Ultra-High-Energy Cosmic Ray (UHECRs) production in Tidal Disruption Events (TDEs) in the light of recent neutrino-TDE associations. We use an isotropically emitting source-propagation model, which has been developed to describe the neutrino production in AT2019dsg, AT2019fdr, and AT2019aalc. These TDEs have strong dust echoes in the infrared range, which are potentially linked with the neutrino production. A mechanism where neutrinos originate from cosmic ray scattering on infrared photons implies cosmic rays in the ultra-high energy range, thus suggesting a natural connection with the observed UHECR. We extrapolate the three TDE associations to a population of neutrino- and UHECR-emitting TDEs, and postulate that these TDEs power the UHECRs. We then infer the source composition, population parameters, and local rates that are needed to describe UHECR data. We find that UHECR data point towards a mix of light to mid-heavy injection isotopes, which could be found, e.g., in oxygen-neon-magnesium white dwarfs, and to a contribution of at least two groups of TDEs with different characteristics, dominated by AT2019aalc-type events. The required local TDE rates of , however, are more indicative of the disruption of main sequence stars. We propose an enhanced efficiency in the acceleration of heavier nuclei that could address this discrepancy. The predicted diffuse neutrino fluxes suggest a population of astrophysical neutrino sources that can be observed by future radio neutrino detection experiments. The derived source parameters are consistent with those expected from the individual neutrino observations.
UTF8gbsn
1 Introduction
Tidal Disruption Events (TDEs) are astrophysical transients originating from stars that are tidally destroyed by a supermassive black hole (SMBH). The accretion of bound stellar materials could power a month to year-long flare (M. J. Rees, 1988; E. S. Phinney, 1989), whose spectrum extends from radio to X-ray frequencies. There have currently been 170 TDEs, TDE candidates and TDE suspects (S. van Velzen et al., 2011; E. Hammerstein et al., 2022; M. Masterson et al., 2024). They have been identified by the Zwicky Transient Facility (ZTF, E. C. Bellm et al., 2019), the Wide-field Infrared Survey Explorer (WISE, E. L. Wright et al., 2010), and X-ray/radio surveys, such as eROSITA (S. Sazonov et al., 2021) and Very Large Array Sky Survey (VLASS, M. Lacy et al., 2020). Remarkably, the correlation study of an 200 TeV IceCube neutrino event IC-191001A and optical transients has revealed that a TDE, AT2019dsg, could be its most likely astrophysical counterpart (R. Stein et al., 2021). The subsequent joint analyses of IceCube neutrino events and electromagnetic signals have discovered more TDE candidates with potential neutrino associations, including AT2019fdr (S. Reusch et al., 2022), AT2019aalc (S. van Velzen et al., 2021), two dust-obscured TDE candidates (N. Jiang et al., 2023), and AT2021lwx (located at a relatively high redshift , C. Yuan et al., 2024a).
The potential correlations between TDEs and astrophysical neutrinos provide the smoking-gun evidence of cosmic ray accelerations since high-energy neutrinos are generated from the interactions between energetic nuclei and ambient radiation/matter backgrounds. However, for TDEs, the sites for particle accelerations and interactions are still unclear. Many models including relativistic jets (X.-Y. Wang et al., 2011; X.-Y. Wang & R.-Y. Liu, 2016; L. Dai & K. Fang, 2017; C. Lunardini & W. Winter, 2017; N. Senno et al., 2017), accretion disks (K. Hayasaki & R. Yamazaki, 2019), wide-angle outflows/hidden winds (K. Fang et al., 2020), and tidal stream interactions (L. Dai et al., 2015; K. Hayasaki & R. Yamazaki, 2019) have been proposed as the origin of the non-thermal electromagnetic (EM) and neutrino emission from TDEs, which could potentially explain these TDE-neutrino coincidences (e.g., R.-Y. Liu et al., 2020; K. Murase et al., 2020; W. Winter & C. Lunardini, 2021; H.-J. Wu et al., 2022; K. Hayasaki, 2021; W. Winter & C. Lunardini, 2023; C. Yuan & W. Winter, 2023).
The non-detection of -rays from these neutrino-coincident TDEs by the Fermi Large Space Telescope (LAT) exerts stringent constraints on the TDE source parameters. In particular, the recent investigation of electromagnetic radiation induced by secondary particles of hadronic (e.g., and ) processes demonstrates that the production of Ultra-High-Energy Cosmic Rays (UHECRs) with energies eV is permitted in an isotropic radiation region of radius cm without exceeding the -ray upper limits (C. Yuan & W. Winter, 2023). Moreover, it was observed that the neutrino-TDE associations are accompanied by strong dust echoes in the infrared wavelength band, and that the arrival times of the coincident neutrinos were close to the peak of the infrared luminosity (which is delayed with respect to the OUV blackbody peak). This led to proposing cosmic ray scattering on infrared photons as a primary neutrino production mechanism, where cosmic-ray primary energies in the UHECR range are needed to exceed the photopion production threshold (e.g., W. Winter & C. Lunardini, 2023; C. Yuan & W. Winter, 2023; C. Yuan et al., 2024a). These conclusions support the idea that TDEs can be promising UHECR sources, and thus offer a possible answer to the long-standing problem of the origin of the observed UHECR flux (G. R. Farrar & A. Gruzinov, 2009; G. R. Farrar & T. Piran, 2014; B. T. Zhang et al., 2017; C. Guépin et al., 2018; D. Biehl et al., 2018b; T. Piran & P. Beniamini, 2023).
The origin of UHECRs is another puzzle in astroparticle physics (R. Alves Batista et al., 2019). It is challenging to associate the measured CRs with a specific source class, (F. Capel & D. J. Mortlock, 2019; A. Abdul Halim et al., 2024a; G. R. Farrar, 2014; V. Shaw et al., 2022). Even with theoretical modeling of promising source classes, it remains unclear which specific source class is responsible for UHECRs, including Gamma-Ray Bursts (GRBs, J. Heinze et al., 2020; D. Boncioli et al., 2019; R.-Y. Liu et al., 2011; K. Murase et al., 2006; F. Samuelsson et al., 2019; M. Vietri, 1995; E. Waxman, 1995) , Active Galactic Nuclei (AGNs, X. Rodrigues et al., 2018; E. Resconi et al., 2017; C. D. Dermer & S. Razzaque, 2010), Starburst galaxies (Romero, G. E. et al., 2018; L. A. Anchordoqui, 2018; L. A. Anchordoqui et al., 1999), and Galaxy clusters (H. Kang et al., 1997; K. Murase et al., 2008; K. Kotera et al., 2009). Sources that produce UHECRs also might emit neutrinos, which contribute to the high-energy diffuse neutrino background. Therefore, the joint consideration of UHECR data and the IceCube diffuse neutrino background would provide valuable insights into the physical environments (e.g., the abundance of intermediate and heavy elements), the acceleration mechanisms (e.g., how the acceleration efficiency depends on the nuclei mass), and redshift distributions of UHECR sources.
In this study, we revisit the potential role of TDEs as sources of UHECRs, motivated by the recent identification of a possible link between three observed TDEs/candidates (AT2019dsg, AT2019fdr, and AT2019aalc) and high-energy neutrino detections by IceCube. While AT2019dsg is well established as a TDE, the classification of AT2019fdr and AT2019aalc remains uncertain, since both occurred in active galaxies where AGN activity can mimick tidal disruption flares; hence, these events are often more broadly referred to as “accretion flares” (S. van Velzen et al., 2021). In addition, the interpretation of AT2019fdr as a superluminous supernova could also account for the coincident neutrino detection (T. Pitik et al., 2022). Based on the models presented in W. Winter & C. Lunardini (2023), we aim to explore the range of parameters that describe TDEs by analyzing various factors, including the type of disrupted stars, the size of the production region, and the maximum energies of CRs. By examining these parameters, we aim to identify scenarios that reproduce the UHECR data and satisfy bounds from the diffuse neutrino flux observations. To achieve this, we develop detailed computational models of cosmic ray emission in TDEs using the NeuCosmA (P. Baerwald et al., 2011; S. Hümmer et al., 2010) software and then model the transport of UHECRs from sources to Earth using the PriNCe code (J. Heinze et al., 2019). By comparing the results with the Akaike Information Criterion, we determine which scenarios best fit current observational data, leading to an improved understanding of TDEs as potential sources of UHECRs.
The paper is structured as follows: In Sec. 2, we develop a model of TDEs using AT2019dsg, AT2019fdr, and AT2019aalc as templates. We identify a set of parameters that will be used in a fit to the data, and determine the ranges of these parameters that are allowed by the observations of these TDEs and TDE candidates. In Sec. 3, we review the current understanding of TDEs as a population of sources, and introduce our scenarios of reference of the UHECR-neutrino-emitting population of TDEs. Using a fitting procedure, we compare these scenarios with the data on the UHECR spectrum and composition from Auger, and present a detailed analysis of the best-fit scenario and how well it compares to the data. In Sec. 5 we elaborate on the question of the consistency with the observed UHECR composition, and discuss a possible enhancement factor that would favor the acceleration of heavier elements in TDEs. In Sec. 6, we discuss the broader consistency of our findings with theoretical and observational expectations, identify areas of tension, and outline directions of future study. Throughout the paper, we use CGS units unless otherwise specified.
2 Modeling of individual TDEs
In this section, we develop and apply a detailed numerical model to simulate the production of UHECRs and neutrinos from TDEs.
2.1 TDE source model
The observed neutrino-TDE associations share certain common properties: relatively high blackbody luminosities in the optical-UV range, strong dust echoes in the infrared range – which are delayed by days with respect to the blackbody peaks – and the neutrinos arriving with comparable delays as well. X-rays were also observed in all these cases. The dust echoes are typically interpreted as re-processed radiation from higher energies, where the delays are indicative of the size of the dust torus surrounding the source. Considering that the neutrino’s time delays are comparable with , and the suggestion that neutrino-emitting TDEs and TDEs with strong dust echoes might be the same population (S. van Velzen et al., 2021), it was proposed that the IR photons might be the main target for the neutrino production (e.g., W. Winter & C. Lunardini, 2023; C. Yuan & W. Winter, 2023) - although a contribution from other wavelengths must be present. This requires that the cosmic ray primary nuclei of mass number and charge number have high enough energies to exceed the photo-pion production threshold. The (kinematical) photopion threshold is given by , where is the nucleon energy, is the target photon energy, is the proton mass, is the pion mass, is the speed of light; here heads-on collisions are assumed at threshold.
In a superposition ansatz, where the energy per nucleon , the threshold condition can be translated into one for the rigidity , where is the nuclear momentum and is the elementary charge:
Thus, for heavy isotopes, where in the stability valley, one needs for target photons in the infrared range, which is independent of the composition to a first approximation. Consequently, the maximal rigidity (which we will use instead of the maximal energy) has to exceed this value, which means that the primary cosmic rays will be in the UHECR energy range. Since the neutrino energy , their interaction will produce neutrinos in excess of about 35 PeV. The emergent discrepancy between the high neutrino energies from UHECR interactions and the observed (mean) neutrino energies below a PeV is a challenge applicable to many source classes for which an UHECR connection is postulated; an exception are sources with strong (kG) magnetic fields such as Gamma-Ray Bursts, where synchrotron cooling of the secondary muons, pions and kaons can break the direct connection between UHECR and neutrino energies, see e.g. W. Winter (2012); we will return to this issue in Sec. 6.
The emergent toy model picture is then a production region of size given by the size of the dust torus, within which the infrared photons produce a quasi-isotropic background photon field. The UHECRs will be accelerated somewhere within this region – such as winds interacting with each other or with the circum-TDE material, off-axis or choked jets – or even in some more compact regions (even though the model probably does not capture all contributions to the neutrino production then).
We adopt the quasi-isotropic emission model in W. Winter & C. Lunardini (2023) (model M-IR) which uses the measured luminosities and temperatures as closely as possible. The interesting photon targets are the blackbody optical-UV emission (OUV, temperature with being the Boltzmann constant), the time-delayed, modelled dust echo in the infrared range (IR, temperature ), and X-rays probably from the innermost regions (temperature ). These photon spectra are modeled by black (gray) body spectra with the individual TDE parameters and time evolutions, see Fig. 3 (black dashed-dotted curves) for two examples. From the (observed) time delays of the IR versus OUV peaks we obtain the size of the dust torus ; within this radius quasi-isotropic IR photons are expected. Depending on the size of the region (inferred from ), on the magnetic field , and on the cosmic-ray energy , this volume can be calorimetric to CRs – i.e., magnetically confined CRs lose their energy by radiation losses – or even optically thick to them.
We assume that cosmic-ray nuclei ZA are accelerated in an acceleration zone which we do not model explicitly, but instead (similar to AGN blazar models) parameterize it by its injection luminosity and maximal rigidity , where . The cosmic-ray injection spectrum (density differential in energy, time and volume) is assumed as
where is defined as the fraction of the energy density injected as element above and is a normalization factor that will be derived from Eq. (3) below. The non-thermal injection luminosity of nuclei with mass number is assumed to follow the mass accretion rate as
where we use as fiducial parameter;111As a consequence, assumes the role of the “baryonic loading”, a parameter commonly used in GRB or AGN models. Note that this formula assumes the injection of a single nuclear species (); later we will use mixed injection compositions whose total luminosities adds up to . we will later discuss its degeneracy with other parameters in Sec. 6. The mass accretion rate is assumed to have a peak value of 222The mass accretion rate is obtained from order-of-magnitude estimations (e.g., C. Yuan & W. Winter, 2023). Noting that during the TDE peak accretion phase, the accretion rate could be super-Eddington, e.g., (L. Dai et al., 2018), where is the Eddington accretion rate that can be related to the Eddington luminosity via using the radiation efficiency . Therefore, we estimate , and adopt the fiducial value for consistency with previous studies. and to trace the observed OUV luminosity time evolution. The minimal energy with , which means that all non-thermal nuclei are picked up (injected) at the same minimal gamma factor (or minimal rigidity) – which is relevant if the mass fraction of the progenitor is to be translated into the non-thermal contribution. Note that the maximal ridigity can be directly converted into the maximal injection energy for a single injection isotope. Later, we will, however, combine different injection isotopes under the assumption of equal , which is the simplest possible assumption. This assumption is obtained if the maximal energy is limited by the size of the accelerator (confinement condition from balancing Lorentz and centrifugal forces). If other processes (such as synchrotron losses or photodisintegration) limit the maximal energy, they will change this relationship, and it may not apply anymore.
In summary, our key parameters, which we vary, are: injection isotope ZA, maximal rigidity and radius of the production region . The studied ranges and are chosen to cover the parameters of the original study for all three TDEs, and to satisfy the constraints from the non-observation of electromagnetic cascades of hadronic origin (C. Yuan & W. Winter, 2023). The maximal rigidity, which earlier was an ad hoc choice beyond the photo-pion threshold of IR photons (W. Winter & C. Lunardini, 2023), is varied in a reasonable range to potentially describe UHECR data, see e.g. J. Heinze et al. (2019). The quasi-thermal photon temperatures, their densities, and their time evolutions are derived from observations of the individual TDEs, as described in detail in W. Winter & C. Lunardini (2023). We fix the magnetic field to be for the sake of simplicity, as in W. Winter & C. Lunardini (2023). A summary of the model’s parameter values is provided in Tab. 1.
Description | Symbol | Value | ||
AT2019dsg | AT2019fdr | AT2019aalc | ||
Input parameters | ||||
Photon temperature [eV] | 0.16 (1) | 0.15 (2) | 0.16 (1) | |
at | 3.4 (3) | 1.2 (2) | 0.9 (1) | |
72 (3) | 56 (2,4) | 172 (4) | ||
Photon luminosity [] | at | at 431 d (1) | at 277 d (1) | at 123 d (1) |
at | (1) | (1) | (1) | |
at | at 17 d (3) | at 609 d (2) | at 495 d (4) | |
Accretion rate factor at peak [] | ||||
Mass of the disrupted star [] | ||||
Dynamical timescale [days] | ||||
Dissipation efficiency [] | ||||
Magnetic field [G] | 0.1 | 0.1 | 0.1 | |
CR maximal rigidity [GV] | ||||
Radius of the production region [cm] | ||||
Characteristics | ||||
Redshift | 0.051 (3) | 0.267 (2) | 0.036 (4) | |
SMBH mass [] | (4) | (4) | (4) | |
Expected number of neutrinos | 0.008–0.76 (3) | 0.007–0.13 (2) | not available | |
Radius of the IR region [cm] | (3) | (2) | (4) |
2.2 UHECR injection in TDEs
We extend the model of W. Winter & C. Lunardini (2023) to nuclei following the methods in D. Boncioli et al. (2017); D. Biehl et al. (2018a, b), which we refer to for details. We simulate the source for different injection isotopes, which disintegrate and interact with the IR, OUV, and X-ray target photons in a test-particle approach, leading to a nuclear cascade in the source. We perform the computation, including the full-time evolution of the target photons, which we also show below; for the main results, we use time-integrated neutrino and UHECR spectra since (at least for UHECRs) the time dependence cannot be measured. We then superimpose the results from different nuclei in the source-propagation chain, which is a possible simplification, since we do not take into account non-linear feedback (e.g., disintegration photons adding to the target).
Our time-dependent solver is based on NeuCosmA, constructing a coupled Partial Differential Equation (PDE) system for all nuclear isotopes considered, protons, neutrons, muons, pions, kaons, and neutrinos, and evolving it with a Crank-Nicolson solver (J. Crank et al., 1947) of the PDE system re-parameterized on a double-log scale in -. The setup of the nuclear system is performed fully automatically with a recursive algorithm following all possible disintegration and beta decay channels beyond a chosen threshold value for the secondary multiplicity (1%) – which acts as a control parameter for the completeness of the system and for the precision of the results. Therefore, the size of the PDE system critically depends on the injection mass number , and the computation time scales correspondingly. Radiation processes taken into account are nuclear disintegration, spontaneous (nuclei) emissions, beta decays, synchrotron cooling losses, photo-pion production, Bethe Heitler pair production losses, escape (discussed below) and interaction losses, whereas we assume a stable production region (), which means that adiabatic losses can be neglected. As a computational example a single run for AT2019fdr with 56Fe injection (largest considered) takes about two hours on a single workgroup server core. Out of 481 possible tabulated isotopes and 39962 tabulated disintegration channels (based on TALYS (A. J. Koning & D. Rochman, 2012)), 218 isotopes and 3838 disintegration channels are automatically selected in that case and evolved over the duration of the TDE. The same computation for proton injection takes only 30 seconds.
corresponds to the evolution of the whole system, and the free-streaming rate is . The shaded region is beyond the assumed maximal injection energy.
For illustration in this section, we follow the original models in W. Winter & C. Lunardini (2023) as closely as possible, which means that we use the same fiducial TDE parameters and also include an outflow with the nuclei may interact with. Here we inject 56Fe with an ad hoc choice (whereas later will be varied). The relevant interaction and energy loss rates for the primary nucleus are shown in Fig. 1 for AT2019dsg. What we see here is that for iron, nuclear disintegration dominates at the highest energies (thin blue curves), the source is optically thick, and the nuclear cascade will efficiently develop. Both the leading disintegration (thin blue) and photo-meson (thick blue, leading to neutrino production) exhibit a triple-hump structure from interactions off (from lowest to highest) the X-ray, OUV, and IR target photons. Note that the solid curves refer to the peak time () of the TDE, and dashed curves refer to the time of the neutrino emission; since the IR dust echo is delayed with respect to the blackbody peak, disintegration (and photo-meson production) of the IR photons dominates at the time of the neutrino emission. Other processes are subleading at first but can become relevant for lighter nuclei over the evolution of the nuclear cascade as disintegration and photo-meson processes lead to nuclei with lower , see D. Biehl et al. (2018a) for details. Our simulation includes this evolution self-consistently in a fully time-dependent computation.
It is an open question how UHECRs escape from the source, where “escape spectra” refer to the cosmic ray emission spectra ejected into the circum-source environment and injected into the UHECR propagation code, which describes transport through cosmic microwave background and cosmic radiation fields. In the present scenario, CRs at lower energies are confined in the production region by turbulent magnetic fields, whereas the electrically neutral neutrons can free-stream out of the production region – unless it is optically thick to interactions. Since UHECRs can escape from the region if the Larmor radius , a suitable parameterization is given by (P. Baerwald et al., 2013)
with the diffusion coefficient in the Bohm-like regime and the free-streaming timescale . The first term ensures that the escape is limited by the free-streaming (ballistic) rate. While a different diffusion coefficient leads to a slightly different result, and also the coherence length of the magnetic field affects the shape of the escape component (see, e.g. J. Becker Tjus et al. (2022)), this approach describes relatively hard escape spectra required from UHECR data, see, e.g., J. Heinze et al. (2019). We show in Fig. 2, left panel, the corresponding cosmic-ray escape spectra group into different mass groups (see figure caption) for AT2019dsg and pure 56Fe injection without transport effects (other than adiabatic losses). It is clear from the figure that the spectral shape below the peak of each component is quickly masked by a different component, which implies that, for hard enough escape spectra, the fit will not be very sensitive to the precise spectral index, and thus to the details of the diffusion model there. The H mass group is driven by protons and neutrons produced in the disintegration, which is suppressed by photo-meson production at high energies. The 56Fe (hypothetical) free-streaming escape spectrum is shown as the dashed curve, which corresponds to the in-source density suppressed by disintegration losses. It is lower than the Fe curve because it represents a single isotope only, and disintegration is very efficient, as can be seen from the shape of the curve and the comparison with low energies, where disintegration does not occur. In fact, the escape is relatively inefficient compared to nuclear disintegration and photo-meson production, as can be seen in Fig. 1.
One could also think about other contributions to UHECR escape. For example, the outflow may transport particles away by advection, which leads to softer escape spectra following the in-source density, see middle panel of Fig. 2. This contribution relies on the presence of the outflow and depends on its velocity. Another contribution would be the final release of all confined nuclei when the system (such as magnetic fields) decay, which we illustrate in the right panel of Fig. 2. This contribution is, however, small because most of the energy of the system has been dissipated otherwise. Since it turns out that the advective contribution cannot describe UHECR data very well, and the final release contribution is small, we further adopt the diffusive contribution only.333For self-consistency, we nevertheless include the effects of the outflow in terms of interactions and an advective escape rate, which is similar to the effect of an adiabatic cooling rate. This leads to slightly smaller predicted neutrino fluences compared to W. Winter & C. Lunardini (2023) (where advective escape/adiabatic cooling was not considered) and to slightly harder proton spectra at low energies (where the protons cannot pile up so much anymore). While we only show the (summed) results over different mass groups in the figures in this section, we transfer all escape spectra of all isotopes in the system to the UHECR transport code at the interface (see below).
Even though we perform a fully time-dependent source simulation, we will, for most of this study, integrate the escaping cosmic-ray spectra over time because time-dependent effects for UHECRs are not measurable. However, the time-dependent escape is theoretically interesting, and the time-dependent spectra can have an effect on the neutrino production time. We, therefore, show in Fig. 3 the time evolution of the diffusive UHECR escape luminosity for the different mass groups (colors) for AT2019dsg (left) and AT2019aalc (right), in comparison to the OUV and IR, and neutrino luminosities (black curves). The absolute luminosities should be interpreted with care because the UHECR luminosities are integrated in the energy range above (observer’s frame), and therefore effects at low energies are not captured. It is interesting that the primary nuclei disintegrate efficiently in the beginning, leading to the time-delayed (by the disintegration loss rate) appearance of lighter nuclei. There also can be a late-time recovery of UHECR escape because the target photon fields decay and the injected UHECRs (following the OUV luminosity) can escape; this contribution is, however, small compared to the escape luminosity closer to the peak. Finally, note that, in the shown examples, the neutrino luminosities (dashed curves) peak close to the actual neutrino observation times better than for the proton model in W. Winter & C. Lunardini (2023). Here, the neutrino production is driven by the protons and neutrons at lower energies (see Fig. 2 and also D. Biehl et al. (2018a)), which are abundantly produced by the nuclear cascade and delayed by the disintegration cooling timescale.
2.3 Expected composition of accelerated cosmic rays from progenitors
In order to accurately replicate UHECR data, a significant portion of the accelerated nuclei must be predominantly from the nitrogen and silicon groups, i.e., have a medium to medium-heavy composition (A. Aab et al., 2017; J. Heinze et al., 2019; P. Plotko et al., 2023). It is most natural to assume that these nuclei originate from the disrupted star and to identify their nuclear composition with the composition of the progenitor. That direct relationship is expected if there are no selection effects at play (e.g., favoring the acceleration of certain nuclear species over others) and if the black hole tidal forces do not change the composition of the stellar matter. Here, we consider several representative nuclear compositions motivated by the different types of disrupted stars. In particular, we examine main sequence stars (MS), ONeMg White Dwarfs (ONeMg-WD), CO White Dwarfs (CO-WD), Wolf-Rayet stars (WR), and red supergiants (RSG) (with composition averaged over their post-main-sequence lifetime), see Tab. 2. MS, RSG, and WR stars have similar mass abundances, primarily consisting of H and He with small amounts of medium and heavy elements. On the other hand, CO and ONeMg-WD stars have mass abundances dominated by medium and heavy elements, making them historically more suitable for UHECRs.
In this study, we chose the injected nuclear composition motivated by stellar evolution (F. LeBlanc, 2010): H, He, C, N, O, Na, Si, and Fe. Due to computation time restrictions and minimal impact on the final results, we couldn’t include all possible elements, so the fractions of elements not selected were added to the closest nuclear isotope. For the composition of the injected CR (), we consider two options. The first option involves a composition of injected CRs matching the mass abundance. This assumption holds under specific conditions, such as having the same minimal Lorentz factor for all particles and a rigidity-dependent maximum energy (). In contrast, the second option treats the injected CR composition as a free parameter and compares it to the mass abundance of a chosen set of stars.
Mass | Star’s stage | White Dwarf | |||
abundance (%) | MS | RSG | WR | CO-WD | ONeMg-WD |
H | 73.90 | 46.46 | 0.01 | 0 | 0 |
He | 24.70 | 36.74 | 98.10 | 0 | 0 |
C | 0.22 | 0.95 | 0.03 | 50 | 0 |
N | 0.01 | 0.30 | 1.33 | 0 | 0 |
O | 0.63 | 2.72 | 0.03 | 50 | 12 |
Na | 0.23 | 0.99 | 0.26 | 0 | 88 |
Si | 0.0 | 0.30 | 0.07 | 0 | 0 |
Fe | 0.12 | 0.52 | 0.14 | 0 | 0 |
2.4 Tidal disruption constraints on SMBH mass from different progenitors
A star is tidally disrupted when its tidal disruption radius exceeds the Schwarzschild radius of the black hole; here is the SMBH mass, is the star’s radius, is the star’s mass and is the gravitational constant. If , the star will cross the event horizon before being torn apart, preventing the formation of a luminous TDE. The maximum SMBH mass for which a star of given mass and radius can be disrupted is given by S. Rosswog et al. (2009):
Description | Symbol | Star’s stage | White Dwarf | |||
MS | RSG | WR | CO-WD | ONeMg-WD | ||
Stellar mass [] | 1.0 | 7.5 | 0.6 | 1.1 | ||
Stellar radius [] | 1.0 | 3.6 | ||||
Upper limit on SMBH mass [] |
Tab. 3 lists the upper limits on SMBH mass for different progenitor types. The SMBHs associated with AT2019dsg, AT2019fdr, and AT2019aalc have estimated masses around , indicating that main sequence stars, red supergiants, and Wolf-Rayet stars can be tidally disrupted in these events. However, white dwarfs (both CO-WD and ONeMg-WD) have much lower values, suggesting that they would likely plunge directly into the black hole before experiencing tidal disruption.
These calculations assume a non-rotating (Schwarzschild) black hole, whereas for of a spinning (Kerr) black hole, the disruption conditions are modified. A rapidly rotating SMBH increases the tidal disruption radius relative to the event horizon, effectively increasing the maximum mass for which disruption can occur. For a maximally spinning Kerr black hole, can be up to 1.5-10 times higher than in the non-spinning case (H. Sponholz, 1994; S. Kobayashi et al., 2004; A. Mummery, 2023). This means that white dwarfs that would otherwise be swallowed as a whole by a non-rotating SMBH could still undergo partial disruption near a rapidly spinning SMBH.
While the direct disruption of white dwarfs in AT2019dsg, AT2019fdr, and AT2019aalc appears highly unlikely, even when considering spinning effects and uncertainties in the SMBH masses associated with these events, we include WD-like compositions in our modeling for comparison. This is motivated by previous studies (B. T. Zhang et al., 2017; D. Biehl et al., 2018b), which indicate that a white dwarf-like composition are necessary to explain the observed cosmic-ray composition at Earth. Although our primary focus is on fitting the composition to the available data, with main-sequence stars as the dominant progenitors, it is important to emphasize this clear tension between the high SMBH masses, which strongly disfavor white dwarf disruptions in these events, and the heavy-element composition inferred from UHECR observation.
2.5 Consistency with individual neutrino-TDE associations
It’s important to validate the consistency of our individual TDE models with the observed data before studying the UHECRs from a population of TDEs. The key observation is the expected number of muon neutrinos for three specific TDEs/candidates: AT2019aalc, AT2019fdr, and AT2019fdsg (see Tab. 1). Our model involves three free parameters: the radius of the production region , the maximum rigidity , and the composition of injected CRs. It’s crucial to identify the allowed regions of these parameters that align with the expected number of muon neutrinos.
We start by calculating the muon neutrino fluence at Earth for each TDE across a grid for both the radius of the production region () and the maximum rigidity (), with the radius ranging from to cm and the maximum rigidity ranging from to GV. We consider pure injection for CRs of a single isotope from the seven possible injection isotopes (H, He, C, N, O, Na, Si, Fe). The next ingredient is effective area , which represents the IceCube detector’s sensitivity to neutrinos across different energy ranges. We utilize the Gamma-Ray Follow-up (GFU) effective area as described in (E. Blaufuss et al., 2019). The predicted number of muon neutrino events by a model is then calculated by combining the predicted fluences with the effective areas over the entire energy range. This calculation can be expressed as:
The next step involves comparing the predicted number of muon neutrinos expected by IceCube. Based on the previous studies, there are certain expectations for the number of muon neutrinos from AT2019dsg and AT2019fdr, but such expectations are not available for AT2019aalc. For instance, the detection of a single high-energy neutrino associated with AT2019dsg suggests a mean expectation in the range of 0.008 to 0.76 muon neutrinos at the 90% confidence level, considering the total TDE population observed by the Zwicky Transient Facility (ZTF) (R. Stein et al., 2021). Similarly, for AT2019fdr, the expected number of muon neutrinos is between 0.007 and 0.13, based on its contribution to the g-band peak energy flux among nuclear transients (S. Reusch et al., 2022). Based on this information, we use the allowed range of 0.007-0.13 muon neutrinos for our study. As indicated in the study by W. Winter & C. Lunardini (2023), the radius of the production region is the main control parameter for the expected number of neutrinos. Consequently, we present our findings graphically against the radius as shown in Fig. 4. For each value of the radius of the production region, we calculate the minimum and maximum numbers of neutrinos across the entire range of maximum rigidity and all injected elements.
The expected number of neutrinos generally depends on four fixed parameters in our model: the size of the production area, the bolometric IR luminosity, and the redshift. AT2019fdr is the farthest from us , but its bolometric IR luminosity is the highest among the three TDEs ( at 227 days post-peak). For AT2019dsg, the luminosity is the lowest ( at 431 days post-peak), but it is closer , which leads AT2019fdr and AT2019dsg to predict a similar number of neutrinos for the same value of the radius. AT2019aalc has the best combination, being the closest and having a high bolometric IR luminosity ( at 123 days post-peak), making AT2019aalc the most efficient neutrino TDE among all three.
In terms of the allowed region for the radius, the expected numbers of neutrinos for AT2019fdr and AT2019dsg, shown as the grey-blue and red curves in Fig. 4, fall within the allowed neutrino region (marked as a grey area on the plot) only at smaller radii. AT2019dsg shows agreement with expected values only at the smallest radii from to cm. AT2019fdr indicates a preference for smaller radii falling within the range of to cm. Since AT2019aalc is the most efficient neutrino TDE, the allowed region is shifted to a bigger radius value from to cm. Besides the restriction from the expected number of muon neutrinos, there is another independent method to estimate upper limits on the production radius involving IR observations. By interpreting the time-delayed IR emission as a dust echo, the size of the production region can be estimated (see Tab. 1 and corresponding references). For all TDEs, the allowed region based on the expected number of neutrinos is inside the one based on IR data (grey arrows on the bottom plot in Fig. 4).
Apart from influencing the expected number of neutrinos, the production radius also significantly affects the CR disintegration within the radiation zone. Fig. 5 shows the escaped spectra of CRs and neutrinos for two distinct radii — and cm — while maintaining the same maximum rigidity ( GV) and CR injection (pure Si). A smaller radius increases the CR interactions with background photons before they can escape, leading to a higher flux of secondary particles at similar energies of the injected primary element. Conversely, a larger radius allows CR survival. As demonstrated by A. Aab et al. (2017) and J. Heinze et al. (2019), UHECR data indicate sources with a peaky quasi-monochromatic spectrum of isotopes transitioning from protons at lower energies to iron at higher energies. For our study, this means that a “contamination” of secondary nuclei in the escaped spectrum from disintegration makes the source unsuitable as UHECR source. This especially affects AT2019dsg, for which fixing the radius to smaller values, as required to produce a high enough neutrino fluence, implies such disintegration. Consequently, we assume that the dsg-like TDEs do not significantly contribute to UHECR and exclude it from our analyses. The following section will focus only on AT2019aalc and AT2019fdr.
3 Modeling of UHECR-neutrino emitting populations of TDEs
In this section, we will discuss the challenges associated with creating a model for a UHECR-neutrino emitting population of TDEs based on two specific models of TDEs: AT2019aalc and AT2019fdr. Our focus will be on numerically simulating the propagation of UHECRs and neutrinos from various options for a population of TDEs, in order to fit the observed spectrum and composition of UHECRs without exceeding the current predictions for the diffuse neutrino flux measured by IceCube.
3.1 Black hole mass function and TDE rates
For each class of TDE (i.e., TDEs involving MS stars, WD, etc.), we use the total cosmological rate of TDEs (number of TDEs per comoving volume per unit time) as a function of , of the form . The redshift dependence reflects the (negative) cosmological evolution of the SMBH and IMBH mass function, as described in, e.g. F. Shankar et al. (2009) (see also N. C. Stone & B. D. Metzger (2016); C. S. Kochanek (2016); C. Lunardini & W. Winter (2017)). The normalization, , depends on the details of the galactic population of stars and of SMBH/IMBH, and in particular on the disruption rate per black hole, , and on the black hole occupation fraction, (the probability that a SMBH or IMBH is located at the center of a host galaxy).
For main sequence star disruptions, a major source of uncertainty is the interval of SMBH mass, , for which tidal disruption is efficient: varying the lower mass cutoff can cause an uncertainty of about an order of magnitude: (see C. Lunardini & W. Winter (2017)), whereas the effect of the upper cutoff, – the mass beyond which a star is swallowed whole and not disrupted — is smaller.
Disruption of white dwarfs must be significantly rarer, due to several factors. One of these is the smaller tidal disruption radius: in Ref. J. H. Krolik & T. Piran (2011) it is reasoned that - if a similar number density of WD and MS stars is assumed - is suppressed by the ratio of the tidal disruption radii: . Further suppression may come from the flattening of the BH mass function at (see discussion in M. MacLeod et al. (2016) and recent results in H. Cho & J.-H. Woo (2024)), and from the possibly low (and uncertain) occupation fraction for IMBH (H. Cho & J.-H. Woo, 2024). Considering the large uncertainties, rates in the range , are realistic, see, e.g. J. H. Krolik & T. Piran (2011); R. V. Shcherbakov et al. (2013); B. T. Zhang et al. (2017). Extreme values, up to are found using recent numerical simulations (A. Tanikawa et al., 2022).
For post-main-sequence stars, rates are typically lower than for MS stars. For example, the rate is estimated at for red giants (D. Syer & A. Ulmer, 1999), and should be orders of magnitude smaller for red supergiants, which are a lot more rare (a specific rate estimate could not be found for these).
We stress that the estimates above are tentative at best and could change significantly as a more detailed description of the tidal disruption process – taking into account elements like binary star systems, partial tidal disruptions, secondary black holes away from the galaxy’s center, black hole spin, etc. – emerge in the future.
3.2 From two to many: Extrapolating the model to the TDE population
Considering the sparseness of observations from only two selected TDEs, applying our model to the entire UHECR-neutrino emitting TDE population is challenging. Rather than attempting to be comprehensive, here we will discuss some plausible, although necessarily simplistic, scenarios. Simplification is necessary to fix the ideas and keep the computation numerically tractable.
A key question here is how to assign values of the fitted parameters and other relevant input parameters (e.g., IR photon temperature and luminosity, etc.) to each member of the UHECR-neutrino emitting TDE population. We are taking a model-independent approach and assuming that each of the AT2019aalc and AT2019fdr represents its own unique group. We will refer to them as the aalc-group and the fdr-group, which together represent the total population. Each group has its unique parameters and local rate (, where aalc or fdr). The local rate of the entire UHECR-neutrino emitting TDE population is given by . The analyzed range and values of these parameters are summarized in Tab. 4
Our analysis works as follows: the TDE population is divided into two groups of identical objects: aalc- and fdr-groups. Each group is described by fixed parameters that are determined by the data on the two selected TDEs and several variable parameters that will be fitted to the observed UHECR data. These parameters are the maximum rigidity (), the radius of the production region (), local rate , and the composition of injected CRs ().
Considering the large number of fitting variables and the limited amount of data, we have to consider different scenarios, which are compared to each other. In particular, we examine the following:
1.
Mono Scenario (Dominance of a single TDE group). In this scenario, we assume that one of the TDE groups is dominant while the other group is suppressed. For example, the aalc-group dominates when . This approach helps us ensure that we don’t overfit the data with a more complex model and allows us to test the scenario where a predominantly single group can explain the observed UHECR data.
2.
Democratic Scenario (Same parameters for all TDE groups). Here, all the fitting parameters are considered the same for the entire TDE population. That represents the situation where the local rates of the groups are democratic (same ), all the disrupted stars belong to the same class (e.g., main sequence stars), and the UHECR production mechanisms are as similar across the population as the current data allow. In this scenario, the production region for the aalc- and fdr-groups overlaps only within a small region around cm, based on constraints from neutrino data for AT2019aalc and AT2019fdr.
3.
Diverse Scenario (Different parameters for all TDE groups). Here, we describe a more realistic situation where the TDE population has substantial diversity. We consider scenarios where the aalc- and -fdr groups have different fitting parameters. This allows for variations in physical conditions and disruption rates across different types of TDEs.
For each scenario, we are considering two options for the composition of the injected CR, as discussed in Sec. 2.3. The first option follows the maximum abundance of a star from the selected statistics. The second option treats the composition of the injected CR as fitting parameters.
Fitting parameters | aalc-group | fdr-group |
[ | Not restricted | |
Injected CR fraction [%] | MS, CO-WD, NeMgO-WD, or Fitted |
3.3 UHECR transport to Earth and description of UHECR data
In our analysis of UHECRs originating from TDEs, we employ the open-source software PriNCe (J. Heinze et al., 2019), designed to numerically solve the one-dimensional transport equations, to model the propagation of CRs in the Universe and compute the cumulative cosmogenic neutrino spectra. We propagate the CRs escaped from the source for each aalc- and fdr-group, considering varying radius, maximal rigidity, and injected composition. We consider a redshift evolution up to for the TDE populations to calculate the cumulative diffuse neutrino fluxes. For the configuration of the calculation, we use the build-in Talys (A. J. Koning et al., 2007) nuclear interaction model, and use the Extragalactic Background Light (EBL) model by R. C. Gilmore et al. (2012).
Our primary goal is to fit the UHECR spectrum (V. Verzi, 2019) and composition (A. Yushkov, 2019) data above , observed by the Pierre Auger Observatory (Auger). In addition, we ensure that the diffuse neutrino flux from the TDE population does not exceed the High-Energy Starting Events (HESE) data from IceCube. For the composition fitting, we attempt to reproduce the observed mean depth of shower maximum density () and its standard deviation (), using the EPOS-LHC air-shower model (T. Pierog et al., 2015)444We also tested the Sibyll 2.3d air-shower model (F. Riehn et al., 2020) and found no significant difference. The results from Sibyll 2.3d air-show model are available at plotkopavlo.com/uhecrsources/tde.. To account for systematic, we include the energy and uncertainties. However, we exclude the uncertainties since they have a minor influence on our analysis.
To evaluate the goodness of the fitting, we adopt the method from P. Plotko et al. (2023) and calculate the value for the corresponding model of the TDE population:
To ensure that the modelled spectrum does not exceed the UHECR fluxes at low energies, we incorporate five data points in the energy range from to . The term is a penalty to ensure that our model does not exceed the HESE diffuse neutrino flux. We consider the values of and , defined as the differences between the model predicted results and the observations, within the uncertainty range of corresponding data set.
Further, to compare different models of the TDE population, we compute the Akaike Information Criterion corrected (, H. Akaike, 1974; K. P. Burnham & D. R. Anderson, 2004; Buchner, J. et al., 2014; A. Rosales et al., 2020) through the given equation:
where is the number of data points, and is the number of parameters of the corresponding model . The model with the smallest value explains the data best among tested models. It is important to highlight that should not be used as a null-hypothesis test (D. R. Anderson et al., 2000; R. Mundry, 2011). In our study, we compare the score of each model against of the model with the smallest value (), calculating the difference () as:
To make it easier to observe the differences between models, we compute in the units of standard deviations following P. Plotko et al. (2023), and use it as an identifier for model selections.
4 Results: fitting TDE population models to UHECR and neutrino Data
4.1 Evaluation of TDE population models
We evaluate various configurations of UHECR-emitting TDE populations to identify models that accurately describe the observed UHECR spectrum and composition distributions without exceeding the constraints of IceCube’s diffuse neutrino flux. The statistical comparison of these models is presented in Tab. 5, including values, degrees of freedom, and AICc values.
As expected, the Diverse Scenario best explains the data among the considered scenarios. In this scenario, the aalc- and fdr-groups have distinct fitted parameters, allowing for a better match to the observed UHECR spectrum and composition. The Democratic Scenario, where all groups share the same parameters, is excluded by more than 8, indicating that uniform fitted parameters across all groups cannot effectively explain the UHECR data. The Mono Scenario, which assumes only one dominant group while excluding the others, is only viable when the aalc-group is dominant, but it is still excluded by . These findings suggest that at least two distinct groups with different fitted parameters are necessary to reproduce the observed data. Hence, we examine the Diverse Scenario more closely in the following subsections.
In cases where the composition of injected CRs is based on the star mass abundances, the performance is worse compared to the Fitted Composition Case, where the composition of each nuclei group is allowed to vary freely. The Diverse Scenario performs the best when the TDEs from the aalc-group disrupt the ONeMg-WD, and the TDEs from the fdr-group disrupt the WR. Although, this case is still excluded by 4.1 compared to the Fitted Composition Case. Based on these findings, we present results only for the Fitted Composition Case of the Diverse Scenario.
. Scenario Composition Group’s local rate Fit quality metrics aalc-group fdr-group aalc-group fdr-group / d.o.f. Mono aalc MS 10 654/29 603 CO-WD 59 84/29 33 5.3 ONeMg-WD 75 192/29 140 Fitted 122 57/27 12 2.8 fdr Fitted 9 261/27 216 Democratic MS 69 37650/29 37599 CO-WD 75 2420/29 2369 ONeMg-WD 69 447/29 396 Fitted 69 248/27 203 Diverse ONeMg-WD WR 47 5 57/24 23 4.3 Fitted Fitted 194 13 38/25 Best-fit case
4.2 Describing the UHECR spectra and composition with TDE population
In this section, we present results for the best-fit case of the UHECR-neutrino emitting TDE population, which is the Diverse Scenario in the Fitted Composition Case. In this scenario, the aalc- and fdr-groups have different parameters, such as the radius of the production region, the maximum rigidity, and the injected CR composition. Fig. 6 illustrates the predictions for UHECR spectrum and composition and diffuse neutrino flux with a comparison to the data for the best-fit model. The obtained fitting parameters for the Diverse Scenario are detailed in Tab. 6.
Our analysis shows that the predicted UHECR spectrum closely aligns with the spectrum observed by the Auger Observatory, with the fitted value of the systematic energy shift of about -9.4%. The aalc-group primarily dominates the UHECR spectrum. Although the fdr-group’s overall contribution is subdominant in spectral fitting, it plays a crucial role in composition fitting, particularly at the highest energies. The model adjustment results in a heavier composition by shifting by approximately 174%. Notably, this adjustment is consistent with recent findings from Auger, e.g., a potential systematic bias toward a heavier composition (A. Abdul Halim et al., 2024b).
In this scenario, the diffuse source neutrino flux from the TDE population does not exceed IceCube’s HESE data. The contribution becomes significant at energies of GeV and higher, peaking at GeV. Unlike the UHECR spectrum, both the aalc- and fdr- groups contribute at a comparable level to the diffuse neutrino background (see the upper right panel of Fig. 6). However, we also observe that the aalc-group produces a more peaked spectrum with a lower maximum energy in contrast to the fdr-group. As a result, the two groups predict different slopes of the neutrino spectrum in the EeV energy range. As can be seen on the upper right panel of Fig. 6), both groups are likely detectable by forthcoming neutrino detectors like IceCube-Gen2 and GRAND200k. Regrettably, the predicted diffuse flux is too low for RNO-G to detect.
Regarding the composition of injected UHECR, the aalc-group emits CRs in both intermediate-mass (e.g., sodium or silicon) and the light (e.g., hydrogen and helium) ranges, whereas the fdr-group mostly contributes the heavy nuclei in the best-fit scenario. This distribution is depicted in Fig. 7 and detailed in Tab. 6. The composition fitting for the aalc-group is similar to ONeMg-WD. Still, it requires additional light elements. The absence of intermediate elements could be attributed to the loss of intermediate elements before they are transferred to the acceleration zone. Furthermore, more irons are needed for the fdr-group to make the predicted composition heavier and better fit the composition data. However, within the 1 AICc region, the injected composition could be pure silicon, which is similar to the ONeMg-WD. One way to achieve an iron-abundant CR injection for both groups is by disrupting a MS star with an enhancement factor for the acceleration of heavier elements, which will be discussed in the Sec. 5.
aalc-group | fdr-group | |
Injected CR fraction [%] | ||
HHe | ||
CNO | ||
NaSi | ||
Fe | ||
38/25 |
In order to match the observed flux of UHECR, a local rate of 200 is required for the aalc-group. This value is typical for TDEs that disrupt MS stars rather than white dwarfs. For the fdr-group, the local rate should be around 10 , which is more indicative for post-main-sequence stars.
Further, Fig. 8 explores the parameter spaces for the aalc- and fdr-groups, and Fig. 9 presents the escaped spectra of CRs and neutrinos for the best-fit parameters. The aalc-group parameters are notably similar to those previously modeled for AT2019aalc (W. Winter & C. Lunardini, 2023), with a maximum rigidity value of GV that is expected for UHECR sources (J. Heinze et al., 2019; A. Aab et al., 2017; P. Plotko et al., 2023). The composition 3 region requires values between and GV to fit the data. Higher values are excluded since, in this case, the escape CRs from TDE are not dominated by a single element group but also have too many secondary light elements, contradicting the data.
On the other hand, the 3 allowed region, based on the spectral fitting, for the aalc-group (marked as the yellow region in the left panel of Fig. 8) is broad and does not significantly impact the fit. The spectrum is less sensitive to composition and does not provide a clear reason for this shape of the 3 spectrum allowed region. The required value of the TDE source radius of the aalc-group is cm, falling within the purple region allowed by the neutrino constraints for AT2019aalc.
Conversely, the fdr-group requires substantially different parameters from previously used in W. Winter & C. Lunardini (2023). In order to fit the composition data, the fdr-group should have a higher maximum rigidity, with the best value of GV. This suggests that the TDEs from the fdr-group could be a powerful UHECR source capable of accelerating iron to energies of GeV. The preference for a high maximum rigidity arises because iron nuclei are required to contribute at energies around GeV, while their presence at lower energies should be negligible. Spectral constraints are not applied to the fdr-group since its overall contribution to the UHECR spectrum is too low to significantly impact the fit.
One may wonder why in Fig. 6 the UHECR flux for the aalc-group is significantly higher than that of the fdr-group (upper left panel), whereas the neutrino fluxes are comparable (upper right panel). As one can see in Fig. 9, fdr-like TDEs produce a (relatively) higher neutrino flux (compare dashed/dashed-dotted curves with solid black curves). The reason is the higher optical thickness due to interactions, driven by the higher luminosities (see Tab. 1) in combination with the much smaller (see Tab. 6).
5 Enhancement of heavier nuclei in the acceleration process
In order to further investigate the possibility of obtaining a heavier CR composition, we introduce an enhancement factor for the acceleration of heavier elements. This option is attractive in combination with the very abundant MS disruptions, which however exhibit a light progentitor composition. Realistically, nucleus selection effects may exist and cause the composition of the accelerated nuclei to differ from that of the disrupted star. One such effect might be partial disruption, which is supported by recent observations (J. J. Somalwar et al., 2023). In this case, the core of the star would remain intact while the matter in the outer layers is disrupted and subsequently accreted onto the central black hole. Ultimately, the accreted material could be accelerated, resulting in the composition of the observed UHECRs. Another possibility is that there might be enhancement mechanisms in the acceleration process that would favor heavier nuclei, thus skewing the UHECR toward heavier composition compared to the original star.
We define an enhancement factor based on D. Caprioli et al. (2017) and A. Hanusch et al. (2019), as D. Caprioli et al. (2017) showed that nonrelativistic collisionless shocks can preferentially accelerate ions based on their mass/charge ratio ( ratios). As a result, these shocks develop a non-thermal tail whose normalization is enhanced as . A. Hanusch et al. (2019) further supports the acceleration of heavier ions through hybrid simulations, which reveal that the injection efficiency of ions into the diffusive shock acceleration process depends strongly on the shock Mach number. Their simulations showed that the injection efficiency for ions with high ratios is enhanced. To generalize the idea, we assume that injection efficiency is proportional to , where is the enhancement factor. Additionally, we assume that thermal particles lose an electron () before being transferred to an acceleration zone to maximize the effect. The net effect is that the abundance of the element in the accelerated stream is scaled by a power of the mass number, leading to new, rescaled mass fractions, :
To ensure that the total mass available for each element (“nuclear budget”) is not exceeded, we calculate the required mass of the injected elements and compare it to the initial mass available in the disrupted star. We will mention if constraints from the nuclear budget require lowering the UHECR luminosity from individual TDE. In Fig. 10, we show the rescaled fractions for our representative stellar compositions, depending on , ranging from (no scaling) to .
In our analysis, we use the Diverse Scenario and focus on the fixed composition of a MS star with various values of for injected CRs. In this scenario, we assume that the aalc- and fdr-groups have different values. For each combination of and , we conduct a fit to determine the minimum value while minimizing over all other fitted parameters. We show as a function of the enhancement factor in Fig. 11. For the aalc-group, the fit requires the value of to be between 1.5 and 2. For the fdr-group, there is a slight preference for values of for the fdr-group, but not statistically significant due to the small contribution of the fdr-group to the UHECR spectrum.
Overall, the Enhancement Factor Case is excluded by more than 8 compared to the Fitted Composition Case due to the nature of the enhancement factor and the assumed fixed composition. The model predicts a light composition for values of the enhancement factor near zero, which contradicts the UHECR composition observations. As the value of increases, the model becomes too efficient at accelerating heavy elements. However, to explain UHECR data, an intermediate-mass group is needed. Different assumptions could help address this issue. For example, there may be two groups of TDEs: one that disrupts white dwarfs with intermediate mass elements and another that disrupts MS stars with an enhancement factor for acceleration of heavier elements. This approach would provide both lighter and heavier elements, potentially improving the fit of the composition data. However, such a model would introduce too many free parameters and could be statistically excluded in favor of the simpler Fitted Composition Case.
Fig. 12 compares the best-fit case to the Auger spectrum and composition, along with the predicted diffuse neutrino flux, for the Enhancement Factor Case. The corresponding parameters of the best-fit case are presented in Tab. 7. While the model’s predictions can explain the UHECR spectrum, the main contribution comes from the iron group, which contradicts the current UHECR composition observation. Additionally, the diffuse neutrino flux from the aalc-group exceeds the HESE limits to a high local rate of 1116 . These two factors result in a poor quality of fit. One potential solution to improve the fit could be using a more complex enhancement factor, such as having iron elements lose more electrons than the silicon group, resulting in a more silicon-dominant composition.
aalc-group | fdr-group | |
1.5 | 1.9 | |
Injected CR fraction [%] | ||
HHe | ||
CNO | ||
NaSi | ||
Fe | ||
104/24 | ||
6 Discussion
Let us now take a broader view of our results and discuss their significance, implications, and limitations.
Choice of parameters, degeneracies.
In this study, we fixed several key parameters, such as the SMBH mass, the dissipation efficiency, the minimum energy, and the spectral index of injected nuclei. We also took into account that certain parameters – including the SMBH mass and the IR, OUV, and X-ray luminosities – are constrained by observations. For instance, we fixed the non-thermal cosmic ray luminosity to be , where the mass accretion rate is taken to be proportional to the Eddington luminosity (and therefore to the SMBH mass). We then derived the local TDE rates that are required to power the observed UHECR flux. These parameters are, however, somewhat degenerate. Indeed, lower can be compensated by a higher local rate or a higher (higher SMBH mass).
Similarly, a higher minimum nuclei energy or a harder injection spectrum lowers the energetic requirement as more non-thermal power is allocated at higher energies, see discussion in W. Winter (2024). Due to this degeneracy, our derived local rates should only be interpreted as indicative. Moreover, degeneracies should be taken into account when comparing with other works; for example,
D. Biehl et al. (2018b) only fit the product of local rate and baryonic loading.
Cosmological TDE rates, connection with jetted TDEs. Despite the existing degeneracies, it is interesting to compare the TDE rates found in this work with those expected from theory and observations. The theoretically expected local rate of all TDEs, (depending on the minimal SMBH mass; see Sec. 3.1) serves as an upper limit for the rate of the sub-class of UHECR-emitting TDEs. Our fit results are consistent with such limit: we found yr-1Gpc-3 for the aalc-group and yr-1Gpc-3 for the fdr-group, which amount to a fraction of the maximum value.
It is noteworthy that our TDE fraction range is roughly compatible with the fraction of jetted TDEs, to (T. Piran & P. Beniamini, 2023), which suggests a possible connection with the existence of jets in TDEs. In the model of W. Winter & C. Lunardini (2023), the jet, which can be completely off-axis, may be the accelerator; the non-thermal UHECRs isotropize in the larger region enclosed by the dust torus, thus making the source effectively a quasi-isotropic emitter. On the other hand, a jet directly pointing towards us, such as in the model of W. Winter & C. Lunardini (2021), is expected to lead to significant electromagnetic signatures (e.g., non-thermal X-rays) which have not been observed in the three cases of neutrino-coincident TDEs but may account for the luminous X-ray afterglows of jetted TDEs, such as the recently reported AT 2022cmc (e.g., I. Andreoni et al., 2022; D. R. Pasham et al., 2023; C. Yuan et al., 2024c, b). Earlier approaches, such as D. Biehl et al. (2018b), also used jetted TDEs for the combined description of UHECR and neutrino data. Note that from the UHECR perspective, there is no conceptual difference between those earlier models and our approach, as UHECRs are deflected by extragalactic magnetic fields and, therefore, do not directly trace the original direction of the jet. In fact, the value of the local (apparent) rate of jetted TDEs that was assumed in D. Biehl et al. (2018b), 0.1 yr-1Gpc-3, corresponds - after taking into account the beaming factor (opening angle from T. Piran & P. Beniamini (2023)) – to a local TDE rate 200 yr-1Gpc-3, which agrees very well with the values we found for the aalc-group.
For neutrinos, however, a jetted model may lead to a higher flux at Earth, due to the contribution of neutrinos from internal shocks inside the jet, which can be very efficient. This additional flux comes from jets pointing in our direction, such as in D. Biehl et al. (2018b).
Type of disrupted stars, nuclear composition. We found that the best-fit composition for the aalc-group is a mix of medium- and light-mass elements. It is overall similar to that of ONeMg-WD (in accordance with B. T. Zhang et al. (2017)), but requires additional light elements. The latter feature might be due to the loss of intermediate elements before they are transferred to the acceleration zone; it could also indicate the existence of two subgroups within the aalc-group, one involving white dwarfs and the other main-sequence stars. However, such a model is beyond the scope of our current analysis with the available data. For the fdr-group, the composition is likely dominated by iron or silicon, within the 1 AICc uncertainty region. Such a composition is also broadly consistent with an ONeMg-WDs.
While the composition analysis favors attributing the bulk of the UHECR flux to the disruption of WDs, several other facts weigh in favor of the disruption of MS stars instead. They are:
The local TDE rates. Our required rates for the aalc-group are well consistent with MS stars, but somewhat in tension with the predicted rates of disruption of WD stars (unless the most extreme values of TDE-WD rates are invoked, see Sec. 3.1).
The estimated black hole masses. In the current understanding of TDEs white dwarfs can only be disrupted by intermediate-mass black holes with typical masses around , while all three observed TDEs involve supermassive black holes around . Based on that, Disruption of the WD is unlikely for AT2019dsg, AT2019fdr, and AT2019aalc. Therefore, for these TDEs, the MS star is a more probable candidate.
There are potential solutions to this tension, where additional physical effects bring the required nuclear composition into consistency with the disruption of MS stars. A possibility is a strong enhancement of heavier elements through the injection processes. A simple, power-law enhancement factor might reproduce the required iron-like composition for the fdr-group (see Sec. 5), whereas for the aalc-group, a more complicated process that favors medium-mass elements would be required. Another idea is that during the disruption of a main-sequence star, central densities, and temperatures can dramatically increase for a very short period, leading to significant nuclear burning. However, there is no clear consensus on whether these conditions are sufficient: J. P. Luminet & B. Pichon (1989) argue in favor, while J. Guillochon et al. (2009) come to a negative answer. Thus, a self-consistent picture has not emerged yet. Such a picture will require a more comprehensive (UHECR-emitting) population model, which would not rely as much on the observed neutrino-TDE coincidences. Indeed, these may not be representative of the whole population; rather, they could be outliers with extraordinarily large neutrino fluences.
During the completion of this work, a new study appeared (P. Milán Veres et al., 2024), revealing a re-brightening of AT2019aalc four years after its discovery. This new flare has characteristics – such as a dust echo and X-ray emission – that are similar to the first one. While the classification of AT2019aalc as TDE still remains unclear, it is speculated that its peculiar properties might be explained by the phenomenology of TDEs in AGN, for example, partial disruption. Spectroscopically, the flare source can be classified as Bowen Fluorescence Flare (BFF, a relatively new class of flaring AGN, L. Makrygianni et al., 2023). Interestingly, another such BFF with a very similar emission spectrum, AT2021loi, has been identified and has been found to be within the (rectangular) 90% confidence region of neutrino event IC-230511. This neutrino arrived 680 (290) days after the first (second) optical peak of AT2021loi.555In fact, an additional neutrino-flare pair was found, involving AT2017bgt and the neutrino IC-200410A, which was, however, a poorly reconstructed cascade event. These observations indicate that, indeed, AT2019aalc might be exceptional and that our aalc-group may correspond to a special type of TDEs in AGN related to BFFs, including AT2019aalc, AT2021loi, and, possibly, AT2017bgt. The spectroscopic results indicate that intermediate to heavy elements, required to describe the UHECR data, are indeed available in the system, even though a direct link to our TDE model requires more information on the origin of the emission lines, which is not yet available. Nevertheless, the limited sample size, classification ambiguities, and potential AGN contamination warrant caution against drawing strong conclusions from these individual events.
Compatibility between observed neutrino energies and predicted neutrino spectra. The direct connection between neutrino energy and CR rigidity (cf., discussion after Eq. (1), applied to nuclei heavier than hydrogen) means that neutrinos are produced at energies following – much higher than the inferred reconstructed neutrino energies in the 100 TeV range for the three neutrino-TDE associations. A quantification of the statistical compatibility between our predicted neutrino spectra (see e.g. Fig. 6, upper right panel) and the energy uncertainty of the IceCube events is nevertheless not possible because uncertainties for the three TDE-associated events have not been published so far, and (especially for muon tracks) the energy estimate is strongly event-dependent and non-Gaussian; it depends e.g. on the uncertainty on the vertex distance and how much energy was lost before reaching the detector (which means it depends on the direction, depth, etc.).
An additional challenge is that the detector response depends on the spectral shape itself. This can be seen, for instance, in the neutrino association of IC-170922A with the flaring AGN blazar TXS 0506+056 for a similar event, where the most probable neutrino energy was given to be 290 TeV: the harder the assumed spectrum in that energy range, the more the probability distribution prefers higher reconstructed neutrino energies (IceCube Collaboration, 2013, see Suppl. Materials, Fig. S2); our predicted spectrum is even much harder in that energy range. A generic attempt to quantify these uncertainties for power law spectra has been made in W. Winter & C. Lunardini (2023) (see e.g. gray-shaded areas in Fig. 10).
This problem has been, in fact, studied in greater detail for AGN blazars in X. Rodrigues et al. (2024), where the predicted neutrino energy follows the maximal proton energy – which is frequently used as free parameter in AGN blazar models: if it is chosen to be in the few-PeV range, it can produce neutrino spectra peaking in the 100 TeV energy range. At these energies, the target photon number is however low, which means that super-Eddington proton luminosities need to be invoked to produce reasonable neutrino event rates, which are incompatible with the standard accretion paradigm in AGN. A higher proton energy can significantly reduce this tension, which leads to the hypothesis that AGN blazar neutrino spectra may also peak at higher energies. X. Rodrigues et al. (2024) have therefore analyzed the detector response for these spectral cases in greater detail (see App. A) and have come to the come to the conclusion that for -like spectra much higher neutrino energies are compatible with the detector response (see Fig. A.2). In fact, the predicted spectra in X. Rodrigues et al. (2024) therefore peak at around the same energies as our TDE predicted spectra. While the theoretical arguments are different here, the conclusions on the detector response apply similarly. However, in F. Oikonomou et al. (2021) the probability that a 20 PeV neutrino is reconstructed with a muon energy as low as 210 TeV was estimated for the individual event IC-190730A associated with an AGN blazar to be around 25%. While such an individual event is consistent with the hypothesis that the spectrum peaks at higher energies, it raises the question if three independent events from three different astrophysical objects with such a spectrum can be found at such low energies, which may be at the level of following their arguments. This line of argumentation would, however, also apply to set of AGN blazars, which raises two more generally applicable possibilities: 1) The observed neutrinos (with the exception of KM3-230213A perhaps, see below) are not coming from a population of sources producing the UHECRs or 2) Some systematics on the detector side leads to an offset in the reconstructed neutrino energies.
Another hint for hard TDE spectra comes from stacking searches, where too soft spectra limit the possibility of neutrino-TDE associations (J. Necker, 2023).
Broader connections: individual extreme UHECR and neutrino events. Regarding the derived UHECR energies, the fdr-group might be able to produce particles with high energies, such as the eV particle observed in 1991 (D. J. Bird et al., 1995) and the Amaterasu particle with an energy of eV in 2021 (R. U. Abbasi et al., 2023). We note that M. Unger & G. R. Farrar (2024) suggests that a transient source may produce the Amaterasu particle, which positions TDEs as a potential candidate. Moreover, N. Bourriche & F. Capel (2024) proposed that the Amaterasu particle could have originated from emitted iron nuclei, which aligns with our results for the composition of the fdr-group. Recently, the detection of a high-energy neutrino event KM3-230213A by KM3NeT-ARCA, with a median neutrino energy of 220 PeV (range 72 PeV to 2.6 EeV at the 90% CL) was reported (S. Aiello et al., 2025), with no clear source association. This energy range is compatible with our diffuse flux computation in Fig. 6, even though our spectrum peaks at slightly lower energies. It is interesting that the interpretation of this event coming from a transient with a lifetime comparable or shorter than the live time of the experiment (335 days) alleviates the tension between IceCube and KM3NeT data (S. W. Li et al., 2025; A. Neronov et al., 2025) – compared to other possible explanations, such as cosmogenic neutrinos.
Fit quality and possible developments. The best value obtained in our analysis was approximately 2.5, which indicates that the fit quality is not ideal. This suggests that while our model captures the broad characteristics of the data, there are likely residuals and complexities in the data that are not fully accounted for. However, attempts to further improve the fit by introducing additional parameters or making the model more complex are unlikely to yield significant benefits. The reason is that the current data may not provide enough additional information to justify an increased model complexity, and such adjustments might only make the theoretical model overly intricate and challenging to interpret. Even in more straightforward cases, such as those discussed by J. Heinze et al. (2019) and A. Aab et al. (2017), achieving a perfect fit has proven challenging.
A potential development of our analysis might be including information on the arrival direction of the observed UHECRs. Considering that this exercise has turned out to be difficult even with simple source models, as indicated by F. Capel & D. J. Mortlock (2019) and A. Abdul Halim et al. (2024a), we feel that applying it to our case of a realistic source model might introduce excessive complication while not adding significant improvements.
Taking a broad look at the situation, we feel that substantial advancements will be possible in the future, when more extensive and detailed multimessenger observations of TDEs become available. These will make it possible to produce more realistic models of TDEs as as a population – or mix of populations – of UHECR and neutrino sources. Those models can then be combined with similarly detailed and robust methods of statistical analysis, to produce state-of-the-art fits of all the available data.
7 Summary and conclusions
For several years, TDEs have been discussed as potential emitters of UHECR and high energy neutrinos. A significant advancement occurred recently, with the proposed associations between neutrinos detected at IceCube and the TDEs/candidates, AT2019dsg, AT2019fdr, and AT2019aalc. All of these events have two common features. The first commonality are strong dust echoes, due to the absorption of optical-UV and X-ray photons by dust and consequent re-emission in the IR band. The peak of the IR flare is observed as delayed relative to the blackbody peak, by the travel time from the central source to the dust region, days. Remarkably, using IR echoes as a selection criterion has led to the identification of AT2019aalc and its associated neutrino. The second commonality is that the coincident neutrinos have all been detected with similar delays of hundreds of days, which place them close to the dust echo peaks. These two facts suggest that the dust echoes may contribute significantly to neutrino production, by providing the target photons for hadron-photon scattering where neutrinos originate. Intriguingly, the primary nuclei must have energies in the UHECR range to exceed the photo-pion production threshold off the infrared targets, therefore the neutrino-TDE associations might be interpreted as indirect evidence for the acceleration of UHECRs in TDEs. In light of these associations, we have revisited the idea that TDEs may power the observed diffuse UHECR flux.
We extended a quasi-isotropic neutrino production model, which was previously proposed to interpret the neutrino-TDE associations, by injecting a variety of nuclear species and allowing the relevant parameters (size of production region and maximal rigidity) to vary. We combined that with a UHECR transport code to predict the cosmic ray composition at Earth. Finally, we performed a fit to the UHECR spectrum and composition data from Auger, and, for each set of fitted parameters, we computed the predicted diffuse neutrino flux at IceCube. It is important to note that the expected local rate of TDEs is determined by fitting the normalization (required emissivity) of the computed cosmic ray spectra to UHECR data, and that predicted diffuse neutrino fluxes are an output (rather than an input) of our calculation.
The main challenges for this work have been: (i) extrapolating from a few neutrino-emitting TDEs to a population of UHECR-emitting TDEs, (ii) reconciling the favored local TDE rate with the nuclear injection composition required by the data, and (iii) ensuring consistency with the individual neutrino-TDE associations. We have identified a best-fit scenario where two populations of TDEs, one being AT2019aalc-like and the other AT2019fdr-like, that describe all relevant UHECR data best while meeting existing constraints. In such a scenario, the aalc-group has an injection composition that consists of a mix of light and intermediate (Na-Si) isotopes, and dominates the UHECR spectrum. In contrast, the fdr-group is characterized by a Fe-like composition and reproduces the highest energy tail of the spectrum. Therefore, our results indicate that at least two different classes of TDEs contribute to the observed UHECR flux. The inferred parameters (radius of production region and maximal rigidity) are found to be in consistency with earlier works. The expected local TDE rates range between 10 and 100 yr-1Gpc-3, where a significant uncertainty is allowed by degeneracies in the parameters. These rates correspond to a fraction of the overall TDE rate, which, interestingly, is compatible with the fraction of jetted TDEs.
From our results, it is not clear what type of stars must be disrupted to power the observed UHECRs. White dwarf disruptions seem to have the right mass composition, but are too rare and associated with lighter black hole masses. Main sequence stars are preferred as progenitors because their local TDE rate is high enough, and the estimated black hole masses and star’s masses of AT2019fdr, and AT2019aalc favor this hypothesis. However, their composition (in terms of the mass fraction) is light. As an option to resolve this discrepancy, we have proposed an enhancement of the injection of heavier elements in the acceleration process, which means that these elements should exhibit larger non-thermal contributions compared to the lighter nuclei. We found that the simplest realization of this idea could make the data compatible with main sequence star disruptions, but a more complicated enhancement mechanism that is most effective for intermediate-mass elements (rather than heavy elements) would be needed to provide a good fit. During completion of this work, evidence that AT2019aalc belongs to a special type of events (which can be identified with Bowen Fluorescence Flares – BFF) was presented, which might be related to (partial) TDEs in AGN. New candidate neutrino associations involving members of this group were identified. It is remarkable that our UHECR data analysis leads to the same conclusion, that the aalc-group might be a special class, thus substantiating the idea of two distinct stellar populations or classes of events further.
In order to associate a specific astrophysical class with UHECRs, reproducing the observed spectrum and composition at Earth is a necessary condition. However, direct source identification remains challenging, as UHECRs are not only significantly deflected by intergalactic magnetic fields but also experience long propagation time delays (R. Mbarek & D. Caprioli, 2025). These delays make it impossible to directly associate UHECRs with transient sources like TDEs. Instead, source and propagation simulations provide a valuable approach to investigating whether certain source classes, such as TDEs, can reproduce the observed UHECR spectrum and composition. At the same time, these simulations help identify key challenges and constraints on the physical conditions required for TDEs to be viable UHECR sources.
We note that the predicted neutrino flux associated with UHECRs inevitably peaks at 10-100 PeV, whereas the observed neutrino energies had much lower inferred energies in the 100 TeV range; see Sec. 6 for a discussion of the compatibility.
On the other hand, such high peak energies suggests that a TDE or a TDE-like AGN flare (C. Yuan et al., 2025) could be the source of the recently observed ultra-high-energy neutrino by KM3NeT-ARCA (S. Aiello et al., 2025). Future radio detection instruments like IceCube-Gen2 (black, M. G. Aartsen et al., 2019), RNO-G (olive, J. Aguilar et al., 2021), and GRAND200k (red, J. Álvarez Muñiz et al., 2019), should be able to robustly test our prediction for the neutrino flux.
In conclusion, we find it overall plausible that TDEs could power the UHECRs in the light of recent neutrino-TDE associations. However, the extrapolation from a few neutrino-emitting TDEs, which may not be archetypic and may be also not very statistically confident in terms of their directionality, to a population of UHECR-emitting TDEs is non-trivial due to limited knowledge and possible selection effects – and some of the parameters, such as the observed neutrino energies, do not support this hypothesis. Independent of that, TDEs are interesting UHECR candidates for different reasons, such as their transient nature, high enough emissivity, and negative source evolution, and our UHECR model and analysis (including the diffuse neutrino flux prediction) remain valid independent of the neutrino associations. Future data will potentially improve the predictions, both on the TDE side and the neutrino side.